OpenStudy (explainitlikeimfive):

Graph the function in the interval from 0 to 2p:

OpenStudy (explainitlikeimfive):

\[y = -4 \cos \frac{ 1 }{ 2 }q\]

OpenStudy (theeric):

I think your questions are a little screwy. Firstly, there's no \(\theta\) in the function, but there's a \(\theta\) on the graph, the horizontal axis. I will assume, however, that \(\theta = q\ \pi\), because I think that's what works out...

OpenStudy (theeric):

\(\Large \theta = \frac{q}{\pi}\), rather..

OpenStudy (theeric):

Well, @ExplainItLikeImFive , what's the amplitude?

OpenStudy (explainitlikeimfive):

Working on it........

OpenStudy (explainitlikeimfive):

aaaaaaaaaaaaaand Im lost in my own work

OpenStudy (explainitlikeimfive):

i dont think it did this right at all

OpenStudy (theeric):

What work did you do?

OpenStudy (explainitlikeimfive):

I tried plugging the equation in, coulndt do it on my own so I gabbed my calculator, and then presumed to error my calculator.

OpenStudy (theeric):

Haha, sounds like a nightmare! So, what's the amplitude? And can that rule out any options?

OpenStudy (explainitlikeimfive):

Yeah I cant figure out the amplitude. But I can rule out B, because thats what I answered and it got marked wrong.

OpenStudy (theeric):

\[y = -4 \cos \frac{ 1 }{ 2 }q \\\ \ \ \ \ \ \ \ \ \uparrow\]

OpenStudy (theeric):

The horizontal stretching or shrinking comes from what you multiply the function by. Otherwise it would be \(1\). Amplitude is the absolute value of that multiplier. Your amplitude is....

OpenStudy (explainitlikeimfive):

4

OpenStudy (theeric):

Right! And that doesn't help, because all of the choices have an amplitude of 4. But there is something. There is no horizontal shift. So, this function would be what when \(x=0\)?

OpenStudy (explainitlikeimfive):

There should be any horizontal movement?

OpenStudy (theeric):

I was actually going to say that - no! Because there is no extra term in the \(cos()\) function! If there was, then yes, but there's not!

OpenStudy (theeric):

The \(\Large\frac{1}{2}\) would cause a stretching, though!

OpenStudy (theeric):

And I think \(q=\theta\), simply. I'm sort of tired! :P

OpenStudy (explainitlikeimfive):

Me to. This is my last question, don'y worry.

OpenStudy (theeric):

Haha, alright, nice!

OpenStudy (explainitlikeimfive):

i'm going with A

OpenStudy (theeric):

Why?

OpenStudy (explainitlikeimfive):

Because its the right answer

OpenStudy (theeric):

Is it really?

OpenStudy (explainitlikeimfive):

Yes

OpenStudy (theeric):

Nuh-uh.

OpenStudy (explainitlikeimfive):

Because its 1Am where I am, therfore it is A.

OpenStudy (explainitlikeimfive):

im totally kidding, its B.

OpenStudy (theeric):

Haha, put \(x=0\) into the function and use a calculator.

OpenStudy (explainitlikeimfive):

B

OpenStudy (theeric):

Really?!

OpenStudy (explainitlikeimfive):

Yes

OpenStudy (theeric):

It's 2am here.... It's either B or D after plugging in \(x=0\), and the \(\Large\frac{1}{2}\) tells you it does cycle normally.

OpenStudy (explainitlikeimfive):

B

OpenStudy (theeric):

The period would be \(4\pi\). Is that so with B?

OpenStudy (explainitlikeimfive):

No

OpenStudy (explainitlikeimfive):

So its D

OpenStudy (theeric):

\[\frac{2\pi}{\frac{1}{2}}=4\pi\] Ahem, final answer?

OpenStudy (explainitlikeimfive):

D

OpenStudy (theeric):

By the way, you've guess all but C. I just found it funny. And I think it's D. Goodnight!

OpenStudy (explainitlikeimfive):

I cannot thank you enough for all the time you spent. I'll let you know what I make on this test, tomorrow, after a good night sleep.

OpenStudy (theeric):

Hopefully you didn't cheat, and you know this stuff! But it seemed like an online "open book" "open internet" test, so what are you going to do? Take care! :)